# spacing=nonfrench

### v1

@@ 6.3 Profiling and Personalization
In this section, we aim to understand if the data collected and shared by the first-party servers of browser assistants is stored and used for profiling of the user. Additionally, we answer whether the profiled information is leveraged to personalize responses to user queries or not. We followed our novel prompting framework described in Section 5.4 to invoke profiling and personalization. Table 4 shows our results. We observe ChatGPT For Google, Copilot, Monica, and Sider
Table 4: Extensions’ Personalization and Profiling Results. * corresponds to logged-out state.
Legend:  signifies profiling or personalization shown in response.  ×  signifies no profiling or personalization shown in response.
Test Prompts Location Age Gender Wealth Interests Profiling Personalization Profiling Personalization (Test 1) (Test 2) (Test 3) (Test 4) (Test 5) (In-context) (In-context) (Out-of-Context) (Out-of-Context) Expected Results Sider Control × × × × × × × × × Search × Browse × × × × × × × × × Summarize × × × × × × × × × Monica Control × × × × × × × × × Search Browse × × × × × × × × × Summarize × × × × × × × × × CFG Control × × × × × × × × × × × × × × × × Search Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × Merlin Control × × × × × × × × × × × × × × × × Search × × × × × × × Browse × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × Max AI Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × Browse × × × × × × × × × Summarize × × × × × × × × × × × × × × × × Perplexity Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × Control* × × × × × × × × × × × × × × × × × Search* × × × × × × × × × × × × × × × × × Browse* × × × × × × × × × × × × × × × × × Summarize* × × × × × × × × × × × × × × × × Harpa Control × × × × × × × × × × × × × × × × × Search × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × Wiseone Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × TinaMind Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × CoPilot Control × × × × × × × × × × × × × × × × × Search Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × ×
@@
to demonstrate expected behaviour for the search scenario where they were able to profile the user based on each leaked attribute regarding location, age, gender, wealth and interests. Additionally, it also personalized user responses both in-context and out-of-context. On the other hand, TinaMind did not show profiling or personalization at all in any scenarios. On a similar note, Perplexity also did not show any profiling, responding to personalization prompt with “I apologize, but I don’t have any prior information about you or your preferences to base personalized recommendations on. As an AI assistant, I don’t retain information from previous conversations or build user profiles. Each interaction starts fresh.”. Moreover, Monica and Sider were observed to not fully profile all leaked attributes about the user in non-search scenarios – only age and interest attributes show consistent profiling in action across scenarios. It is interesting to see that the profiled information transcend beyond context restrictions discussed in Sections 6.1 for Monica and Sider, suggesting that the browser assistant is likely maintaining a profile for the user on server-side to be able to personalize its responses in-context as well as out-of-context in all cases. Surprisingly, no extensions except Monica and Sider, show any profiling or personalization for the contorl scenario. The fact that these two show a positive response in control, suggests that these browser assistants might be associating even the asked questions to be reflective of the corresponding leaked profile.
@@
Harpa demonstrates profiling for all 5 attributes for search and in-context personalization. However, it fails to retain the profile out-of-context. Merlin and MaxAI demonstrate similar profiling behaviour for interest attribute across both search and browse scenarios, resulting in in-context personalization that doesn’t similarly translate to out-of-context. Behaviour of Merlin and MaxAI shows unpredictability – for instance, Merlin search shows profiling at each step of the test prompting. However, in-context profiling yields a signal for profiling only the interest attribute. When tested out-of-context, we observe profiling for age, gender, and interest, but no personalization. We believe the same reasoning could apply here as discussed previously – dependency on the value temperature parameter used by Merlin. Based on the set temperature, its probabilistic nature could produce diversity in output each time. Many extensions result in the output “I’m sorry, but I cannot answer those questions with "YES" or "NO" as I do not have any information about you.”. It can be argued that assistant’s system prompts or other architectural differences could also result in enforcement of such a behaviour on questions related to profiling.
@@
7 Conclusion
In this paper, we systematically evaluated the 10 most popular GenAI-based browser assistants to investigate their architectural design, privacy risks, and profiling and personalization capabilities. Our findings indicate that: (1) 90% of the browser assistants rely on server-side processing, contacting third-party LLM models from their first-party servers by transmitting user queries and metadata for response generation. Notably, two of the extensions (Sider and Copilot) are auto-triggered even when users perform a query on the browser search engine. 60% of the GenAI browser assistants isolate the contexts of each query. However, ChatGPT for Google and Wiseone maintain context across page navigations. (2) Two browser assistants (Harpa and Copilot) collect the full DOMs of user-visited pages, while others gather varying levels of private data from webpage content, such as medical records, student’s academic records, or query-based Google’s search results. Additionally, Harpa and MaxAI share page locations and referrers with third-party tracking services. Merlin was discovered to be the only assistant that recorded the contents of web forms and managed to collect and share SSN as part of a webform. Moreover, some assistants, such as Sider, Merlin, and TinaMind, share chat identifiers and raw queries with Google Analytics, enabling potential tracking and retargeting on external platforms. (3) GenAI browser assistants demonstrate varying degrees of profiling and personalization based on user attributes such as location, age, gender, wealth, and interests. Browser assistants such as ChatGPT for Google, Copilot, Monica, and Sider perform extensive profiling across all five attributes, enabling both in-context and out-of-context personalization when search functionality is used. However, others, such as Perplexity and TinaMind, do not show strong signals of profiling or personalization. Overall, we observed Perplexity to be the most privacy-friendly while extensions such as Harpa, MaxAI, and Merlin were amongst the least.
As highlighted in our work, capabilities of GenAI assistants are tremendous in terms of the granular level of access that these models have to user’s personal or sensitive data. This capability underscores the urgent need for robust regulatory frameworks. Collection and sharing of medical records and student’s academic records by different browser assistants demonstrated non-compliance with the regulations such as HIPAA [14] and FERPA [34]. Browser assistants can easily incorporate a check to decide whether or not the webpage contains PII or sensitive information about the user via domain-based classification. Moreover, they could also adopt Perplexity’s approach to perform a server-side fetch of the webpage based on the URL to avoid collecting any personal data about the user or obtain explicit consent from the user on each webpage similar to Wiseone. In absence of any regulatory enforcement, most of these assistants do not have any mechanisms in-place to evaluate their privacy risks. We recommend policymakers to adopt a bottom-up approach to regulate GenAI assistants as they increasingly influence the future of web browsing and search engines. Privacy must be embedded into these systems by design, and developers must actively mitigate risks of data leakage, profiling, and unauthorized sharing.
As advancements in GenAI shape the future of web browsing, some envision an in-browser model powering an intelligent yet private web browser that can automate different tasks to simplify day-to-day browsing needs of a user. While others anticipate a paradigm shift [44] where traditional browsers and search engines may become obsolete. These transformations are currently being driven by advances in artificial general intelligence (AGI) that aims to automate workflows by integrating different specialized agents that work together to complete user’s task such as booking a dinner reservation in an end-to-end manner. Google is building towards the former vision and has already begun origin trials for Built-in AI [11] to allow developers of websites and webapps to leverage Gemini Nano through several APIs. On the other hand, Microsoft’s Recall [1] recently brought GenAI to user’s computer allowing them to record their screen and integrate recordings with GenAI locally. More recent developments such as Window.ai [2] and Screenpipe [18], have also been in similar directions as Recall, inspiring user-centric designs for managing GenAI interactions directly within the user’s device, reducing reliance on server-side processing and ensuring user privacy. As GenAI browser assistants increasingly shape this landscape, it is important that privacy considerations form the foundation of their design and deployment. Future works should delve deeper into improving transparency mechanisms for GenAI-based browser assistants. Collaboration between developers, researchers, and policymakers will be critical to ensure these tools empower users while safeguarding their rights and privacy. We believe our work will lay a foundational ground to inspire future privacy work in this space.
@@ 7 Conclusion
Our research involves visiting private online spaces such as personal health portal, social media platform, personal email, etc. to understand the risks posed by using GenAI browser assistants in such spaces. We did not recruit any participants for our study and used our own personal accounts to perform the experiments. We reached out to the Institutional Review Board (IRB), but they granted an exemption to our research since it involved the use of personal accounts and therefore did not fall under the definition of human subjects.
Our research involved creating and using accounts with various browser assistant providers to evaluate their services during web browsing experiments. To ensure ethical practices, we utilized temporary email accounts, created no more than one account per assistant per day, upper bounded to 12 accounts, and refrained from excessive account creation. Most of the experiments were conducted using the free token credits provided at sign-up, but to ensure consistency in the features tested across different assistants, we also purchased a monthly paid subscription for some extensions.
Although this study did not involve human subjects, we adhered to ethical principles aligned with those outlined in The Menlo Report. Specifically, we prioritized beneficence by minimizing potential harm, respected the autonomy and integrity of organizations, ensured fairness in the evaluation of different assistants, and remained compliant with applicable laws and public interests.

### v2

@@ 6.3 Profiling and Personalization
In this section, we aim to understand if the data collected and shared by the first-party servers of browser assistants is stored and used for profiling of the user. Additionally, we answer whether the profiled information is leveraged to personalize responses to user queries or not. We followed our novel prompting framework described in Section 5.4 to invoke profiling and personalization. We performed each scenario 15 times. Each cell in Table 4 shows the majority response received across 15 repetitions. A binomial test confirms that this majority outcome (Yes/No) occurs statistically significantly more frequently than expected by random chance with 95% confidence (i.e., p-value $<$ 0.05).
@@
We observe all assistants, except Perplexity and TinaMind to demonstrate in-context profiling and personalization for the search scenario based on almost all leaked attributes regarding location, age, gender, wealth and interests. This is likely because assistants can easily remember the explicitly leaked information through train prompts in the search scenario by keeping the information within their context windows. This is evidently validated through profiling and personalization prompts asked during the test phase, when they were observed to still retain the context. Additionally, Sider, Monica, ChatGPT for Google, and CoPilot also personalized user responses out-of-context (i.e., across browser tabs). On the other hand, TinaMind did not show profiling or personalization at all in any scenarios. On a similar note, Perplexity also did not show any profiling, responding to the personalization prompt with “I apologize, but I don’t have any prior information about you or your preferences to base personalized recommendations on. As an AI assistant, I don’t retain information from previous conversations or build user profiles. Each interaction starts fresh.”. Moreover, although Monica and Sider show the highest profiling and personalization as compared to all other assistants, they were observed to not fully profile all leaked attributes about the user in non-search scenarios – only age and interest attributes show consistent profiling in action across scenarios. It is interesting to see that for these attributes in Monica and Sider, the profiled information transcend beyond context restrictions discussed in Sections 6.1. This suggests that the browser assistant is likely maintaining a profile for the user on server-side to be able to personalize its responses in-context as well as out-of-context in all cases. Surprisingly, no extensions except Monica and Sider, show any profiling
Table 4: Profiling and Personalization Results. * corresponds to logged-out state. Each cells depicts the majority outcome from 15 repetitions of prompting experiments – Yes or No represented in the form of a tick or a cross, respectively. A binomial test based statistically significant results at 95% confidence are colored ( or ). Statistically insignificant outcomes are depicted in black symbols ( or ). Legend:  signifies statistically significant profiling or personalization shown in response.  signifies statistically significantly no profiling or personalization in response.
Test Prompts Location Age Gender Wealth Interests Profiling Personalization Profiling Personalization
(Test 1) (Test 2) (Test 3) (Test 4) (Test 5) (In-context) (In-context) (Out-of-Context) (Out-of-Context)
Expected Results
Sider Control
Search
Browse
Summarize
Monica Control
Search
Browse
Summarize
CFG Control
Search
Browse
Summarize
Merlin Control
Search
Browse
Summarize
Max AI Control
Search
Browse
Summarize
Perplexity Control
Search
Browse
Summarize
Control*
Search*
Browse*
Summarize*
Harpa Control
Search
Browse
Summarize
TinaMind Control
Search
Browse
Summarize
CoPilot Control
Search
Browse
Summarize
or personalization for the control scenario. The fact that these two show a positive response in control, suggests that these browser assistants might be associating even the asked questions to be reflective of the corresponding leaked attribute.
@@
Harpa demonstrates profiling for all 5 attributes for search and in-context personalization. However, it fails to retain the profile out-of-context. It weirdly shows affinity towards inferring and remembering gender attribute during the browse scenario – both in-context as well as out-of-context. Merlin and MaxAI demonstrate similar profiling behaviour for interest attribute across both search and summarize scenarios, but this does not consistently translate to out-of-context scenarios. Behaviour of Merlin and MaxAI also show unpredictability – for instance, Merlin search shows profiling at each step of the test prompting. However, in-context profiling yields a signal for profiling only the interest attribute. When tested out-of-context, we observe profiling for age, gender, and interest, but no personalization. We believe the same reasoning could apply here as discussed previously – dependency on the value of temperature parameter used by Merlin. Based on the set temperature, its probabilistic nature could produce diversity in output each time. Many extensions result in the output “I’m sorry, but I cannot answer those questions with "YES" or "NO" as I do not have any information about you.”. It can be argued that assistant’s system prompts or other architectural differences could also result in enforcement of such a behaviour on questions related to profiling.
To further understand the types of user data retained by these assistants, we leveraged responses to prompts used in the experiments described in Section 6.2.1. We particularly tested for private spaces that are vulnerable to user’s personal or sensitive data. While asking prompts listed in Table 6 after summarization of the webpage, we specifically asked appropriate questions related to user attributes. Particularly, we tested for name, email, height, weight, dating preferences, profile details, personal chats, courses and grades, card details, and viewing interests (highlighted in red in Table 6. We observe that webpage data collected during the summarization phase was retained by the assistants to produce accurate responses in-context as well as out-of-context in non-trivial number of cases. Perplexity and Copilot did not retain any data. Assistants other than these two were able to retain data accurately for some of the scenarios, but missed details in the others. Name, email, height, weight, profile details, academic performance, and personal chats were almost always accurately retained in the follow-up questions. Most of the other details were either partially retained or had missing details from the responses as highlighted in Table 3.
@@
7 Concluding Remarks
In this paper, we presented a novel framework to systematically audit nine popular GenAI browser assistants. We found that most assistants rely on server-side processing, with some (e.g., Sider, Copilot) auto-triggered on search queries, and a few maintain context across browsing sessions and tabs. Several assistants collect sensitive information such as full DOMs, form inputs, and medical data and shares it with third parties. We also found evidence that several assistants infer profiles and uses them for personalization.
Our work shows that GenAI browser assistants offer powerful capabilities but pose significant privacy risks due to their access to personal and sensitive user data. Addressing these risks require efforts across the ecosystem.
Assistant Developers: Assistant developers are central to mitigating privacy risks. Browser assistants can disable data collection on sensitive websites based on domain/URL or keyword lists. Assistants can also prompt users for runtime permission on sensitive websites before collecting, sharing, or processing data.
Extension Stores: Extension stores play a key gatekeeping role. Like mobile app stores, extension stores should adopt stricter review processes for GenAI-powered extensions. GenAI-powered extensions can be required to declare their functionality under pre-defined categories (e.g., personalization, automation) and disclose associated privacy risks to users prior to installation.
Browsers: Browsers can help reduce the risk by offering on-device LLMs for use by browser assistants. While on-device models may not yet match the capabilities of more powerful server-side models, they would eliminate the risk of remote data sharing and processing. Browsers can also define granular permissions to mediate the collection, sharing, and processing of personal/sensitive information by GenAI browser assistants.
Regulators: Regulators such as the Federal Trade Commission (FTC) and European Commission can play an important role in shaping how GenAI browser assistants handle personal and sensitive data [13, 14, 9]. New guidelines on the collection and use of such data for profiling or personalization can help set expectations and define responsibilities for developers, extension stores, browser vendors, and AI model providers. Stronger enforcement of existing privacy laws (e.g., HIPAA, FERPA) can also help mitigate the risks posed by GenAI browser assistants.
Users: End users need to be made aware of the privacy risks of GenAI browser assistants. Transparency initiatives, such as standardized AI usage and risk labels, can help users make informed decisions.
Our work presents the first systematic audit of GenAI browser assistants; however, significant research questions remain unstudied, especially as the ecosystem surrounding GenAI and its browser-based assistants continue to evolve. GenAI is expected to be a key part of the future of web browsing. Our current study focuses on search-integrated Chrome browser assistants. Future audits should expand to other browsers, especially new GenAI-focused browser projects (e.g., Comet [2], Dia [34], Mariner [11]), that are expected to tightly integrate GenAI into the core browsing experience with a shift towards agentic browsing.
@@ 7 Concluding Remarks
We designed and conducted this study with close attention to ethics, following the Menlo Report’s principles of Respect for Persons, Beneficence, Justice, and Respect for Law and Public Interest. All experiments were performed on accounts controlled by the primary researcher (either the researcher’s personal accounts or temporary ones created for this study), with no external individual’s data. The Institutional Review Board (IRB) reviewed our protocol and determined it qualified for an exemption from human-subjects research requirements, as it does not involve any identifiable private information from individuals outside the research team. Nevertheless, an IRB exemption is not a substitute for ethical conduct; we have proactively implemented measures to uphold the highest ethical standards.
Respect for Persons: We took measures to respect individual privacy and autonomy. No other persons were involved at all – the only human subject was the researcher, who consented to their own participation. We prevented any leakage of other users’ data while interacting with private accounts. For each of the 10 private spaces, the researcher used either a personal or new account containing no data from other individuals. We viewed this as a deontological duty to avoid infringing on anyone’s privacy, and consequentially it ensured no harm could come to others from our experiments. More specifically, in private spaces such as Netflix, Tinder, PornHub, IRS and Chase, we did not browse any page containing information about anyone other than that of the primary researcher. In the case of Health Portal, we were logged-in to the patient account of the primary researcher and the browsed content consisted of health information only related to the primary researcher. It did contain the name of the health providers, which is nevertheless public information. For Canvas, the researcher was logged-in as a student and did not browse any page listing information about other students. The only individual other than the researcher that was listed on the browsed course page was the instructor of the course, which also constitutes public information. Slack was used in a private workspace with only the members of the researcher team present. Through these precautions, we maintain the autonomy and privacy of all individuals by ensuring that no one outside the research team was observed or affected in any way.
Beneficence: We carefully weighed the potential benefits of the research against any possible harms, aiming to maximize benefits and minimize harm. The primary benefit is an improved understanding of security and privacy behaviors in emerging GenAI-powered browser assistants, which can inform better protections for users and improved practices by extension developers. The potential harms were primarily privacy-related. By design, these risks were limited to the researcher alone – no other person’s data could be collected or exposed. Only in scenarios where the researcher’s account might incidentally display others’ content (for e.g., an email sender’s name in Gmail), we minimized the harm by using a Gmail account where all the emails were only exchanged with another team member with their consent. Following this, we avoided any incidental exposures from our data, eliminating risk to anyone’s privacy. From a consequentialist perspective, the residual risk—essentially zero for anyone but the researchers—was far outweighed by the knowledge gains and public benefits of our study. From a deontological standpoint, we embraced the principle of “do no harm” by designing the study to preclude harm to others from the outset.
Justice: This principle demands a fair distribution of research burdens and benefits. In our study, no outside individuals were asked to bear any burden or risk – every experiment was performed by members of the research team using their own accounts. Thus, no one was inconvenienced or put at risk. Meanwhile, the insights from this work will benefit a broad extension user and developer community rather than a select few. Researchers within the team voluntarily assumed the minimal burden, while the benefits are widely shared, which aligns with both the consequentialist and the deontological notions of fairness. Our approach maximizes overall good without exploiting anyone, and it treats all individuals with equal respect by not involving them without consent.
Respect for Law and Public Interest: Throughout the project, we complied with all applicable laws, regulations, and (to the best of our knowledge) platform terms of service. All interactions with the private services were through the researchers’ legitimate user credentials; we did not circumvent access controls or engage in any activity clearly prohibited by these platforms. When we employed automated scripts or data collection tools on our own accounts, we ensured this fell under acceptable use. Additionally, to prevent unnecessary resource usage, we employed temporary email addresses and limited account creation strictly to no more than one account per assistant per day, and capped the total number of accounts at 15 (one per experimental repetition). Most experiments utilized free token credits provided upon sign-up; however, to ensure fairness and consistency in comparing features across different browser assistants, we also purchased monthly paid subscriptions wherever necessary. We concluded that our methodology stayed within legal and ethical bounds, serving the greater good without causing undue risk. This approach reflects respect for rules and authority (deontological compliance) as well as outcome-oriented reasoning (consequentialist promotion of societal benefit). Finally, our institutional IRB’s exemption confirmation underscores that the study does not involve human subjects under regulatory definitions. We stress that this IRB outcome does not absolve us of ethical responsibility; throughout the project we maintained rigorous ethical safeguards as if full IRB oversight were in place. In summary, by adhering to the Menlo Report principles through both principle-driven and outcome-driven reasoning, we ensured that our study was conducted in an ethical manner that protects individual rights and welfare while serving the broader interests of society.
Open Science
To foster future research, we open-source our crawling and analysis framework at https://doi.org/10.5281/zenodo.15618960.